Circular No. 7663

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dd IN UGC 11579 Pierre Antonini, Bedoin, France, reports his discovery of a supernova

(mag ∼ 17) on unfiltered CCD exposures taken with a 0.30-m reflector on July 13.07 and 13.88 UT. Confirming images were obtained by J. Lecacheux at the Pic du Midi 1-m reflector on July 14.0. The new object is located at α = 20^{h}30^{m}54^{s}.08, δ = −0^{o}39′26′′.9 (equinox 2000.0), which is 15′′ west and 25′′ south of the nucleus of UGC 11579. Images taken by Antonini on June 22.1 and 26.1 show nothing at this position (limiting mag ∼ 17.5), and the object does not appear on a Digital Sky Survey image taken on 1953 Aug. 16.29 (limiting mag ∼ 21).

N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports: “K. Olsen and S. Thomas observed SN 2001dd with the Blanco 4-m telescope (+ facility spectrograph) at CTIO. The spectrophotometry (range 310–860 nm, resolution 1000) shows that this object is a type-Ia supernova, 7 days after maximum light. Besides strong interstellar absorption lines of the Na I doublet (0.06 nm at 589.0 nm; 0.05 nm at 589.6 nm) due to the Galaxy at heliocentric velocity of −80 km/s, the doublet lines appear at heliocentric velocity 5760 ± 30 km/s with equivalent widths of 0.07 and 0.025 nm (± 0.015 nm), consistent with the published radial velocity of the galaxy (5902 km/s; Haynes et al. 1999, A.J. 117, 2039). SN 2001dd is expected to have significant extinction due to the host galaxy.”

SUPERNOVA 2001cz IN NGC 4679 A. Pastorello, G. Altavilla, S. Benetti, E. Cappellaro, and M. Turatto,

Osservatorio Astronomico di Padova, report: “A fully reduced CCD spectrum (range 310.0–910.0 nm, resolution 1.0 nm), obtained on July 12.01 UT with the Danish 1.54-m telescope (+ DFOSC), shows SN 2001cz (cf. IAUC 7657) to be a type-Ia supernova at maximum light (± 2 days). The spectrum consists of a blue continuum with superimposed P-Cyg lines of Ca II (H and K), S II, and Si II ions. The expansion velocity deduced from minimum of the Si II 635.5-nm doublet is 10 300 km/s (a recession velocity for the parent galaxy of 4612 km/s was obtained from NED). Moreover, the spectrum shows deep interstellar Na I D Galactic (equivalent width ∼ 0.04 nm) and extragalactic (EW ∼ 0.11 nm) absorptions.”

SUPERNOVA 2001bs IN UGC 10018 Corrigendum. On IAUC 7631, line 1, for Beckman read Beckmann

2001 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green